OpenStudy (anonymous):

1. f(x)=x^2; g(x)=1/√x a. f+g b.f-g c.f*g d.f/g e.g/f

OpenStudy (anonymous):

what do you want to find up?

OpenStudy (anonymous):

Define the following functions

OpenStudy (anonymous):

you meant domain?

OpenStudy (anonymous):

i dont know.. that was the only instruction given sir..

terenzreignz (terenzreignz):

\[\Large f(x)=x^2\]\[\Large g(x)=\frac1{\sqrt x}\]

terenzreignz (terenzreignz):

This stuff is actually pretty straightforward. I'll do the first one to give you an idea... f+g Just outright add the two functions, no biggie XD \[\Large f+g = x^2 +\frac1{\sqrt x}\]

OpenStudy (anonymous):

nothing will be add?

terenzreignz (terenzreignz):

There's no way to simplify this, as far as I know.

OpenStudy (anonymous):

same goes with b?

terenzreignz (terenzreignz):

Same goes with all of them, actually, just perform the indicated operation ^^

terenzreignz (terenzreignz):

What's your answer to b?

OpenStudy (anonymous):

x^2+1/√x

terenzreignz (terenzreignz):

That's the answer to a. But what about b?

OpenStudy (anonymous):

x^2-1/√x ? xD

terenzreignz (terenzreignz):

\[\Large f-g = x^2 - \frac1{\sqrt x }\] correct :P

terenzreignz (terenzreignz):

c ?

OpenStudy (anonymous):

x^2 * 1/√x ? xD

terenzreignz (terenzreignz):

\[\Large x^2 \cdot \frac1{\sqrt x}\] But this can be simplified into just one fraction, no? In fact, it doesn't even have to be a fraction :D

OpenStudy (anonymous):

uhh?

OpenStudy (anonymous):

how come?

OpenStudy (anonymous):

oohh.. i see,.. it this right?